# MegaLoc: One Retrieval to Place Them All

### v2

@@ Abstract
Retrieving images from the same location as a given query is an important component of multiple computer vision tasks, like Visual Place Recognition, Landmark Retrieval, Visual Localization, 3D reconstruction, and SLAM. However, existing solutions are built to specifically work for one of these tasks, and are known to fail when the requirements slightly change or when they meet out-of-distribution data. In this paper we combine a variety of existing methods, training techniques, and datasets to train a retrieval model, called MegaLoc, that is performant on multiple tasks. We find that MegaLoc (1) achieves state of the art on a large number of Visual Place Recognition datasets, (2) impressive results on common Landmark Retrieval datasets, and (3) sets a new state of the art for Visual Localization on the LaMAR datasets, where we only changed the retrieval method to the existing localization pipeline. The code for MegaLoc is available at https://github.com/gmberton/MegaLoc
@@ 1 Introduction
Why does it matter?”, you may ask. Imagine you are doing 3D reconstruction, where image retrieval is a fundamental component, on a collection of diverse scenes (e.g. to create datasets like MegaDepth [18], MegaScenes [37], or for the evergreen Image Matching Challenge [6]). In some cases there would be small scenes (e.g. reconstruction of a fountain), requiring a retrieval model that is able to retrieve nearby images (few meters away), which is something VPR models excel at, but LR models underperform (see [8] Tab. 14). In other cases however, the scene might be large (e.g. a big landmark like a church), with images hundreds of meters away: while LR models are designed for this, VPR models achieve poor results in this situations (see Sec. 3). Given these considerations, we note how neither VPR nor LR provide models for the diverse cases of 3D reconstructions, creating a gap in literature that is filled by MegaLoc. As another example where a model like MegaLoc is necessary, one can think of Visual Place Recognition (which is also the first step for Visual Localization), where models are evaluated by using a 25 meters threshold (and queries in popular datasets always have at least one positive within 25 meters). However, in the real world the nearest image to a given query might be 100 meters away, and while ideally we would still want to retrieve it, a VPR model is unlikely to work in such case, as it has been trained to ignore anything further away from the camera.
In this paper we demonstrate that, by leveraging a diverse set of data sources and best practices from LR, VPR and VL, we obtain a single image retrieval model that works well across all these tasks. Our model is called MegaLoc and it is released at https://github.com/gmberton/MegaLoc
@@ 2 Method
The core idea of this paper is to fuse data from multiple datasets, and train a single model. We use five datasets containing both outdoor and indoor images and catering to different image localization tasks: GSV-Cities [1], Mapillary Street-Level Sequences (MSLS) [39], MegaScenes [37], ScanNet [13] and San Francisco eXtra Large (SF-XL) [7]. At each training iteration, we extract six-sub batches of data, one for each dataset (except SF-XL, from which two sub-batches are sampled) and use a multi-similarity loss [38] computed over each sub-batch. Each sub-batch is made of 128 images, containing 4 images (called quadruplets) from 32 different places/classes. Given that these datasets have diverse format, they require different sampling techniques. In the following paragraphs we explain how data is sampled from each dataset.
@@ 3.1 Implementation details
consists of a DINO-v2-base backbone [21] followed by a SALAD [16] aggregation layer, which has shown state-of-the-art performances over multiple VPR datasets [16, 33]. The SALAD layer is computed with 64 clusters, 256 channels per cluster, a global token of 256 and an MLP dimension of 512. The SALAD layer is followed by a linear projection (from a dimension of 16640 to 8448) and an L2 normalization.
@@
is achieved using PyTorch [23], by ensuring that the computational graph for each loss stays in memory as little as possible. In practice (in the code), instead of adding the computational graph for each loss into a single giant graph, we compute each loss and perform the $backward()$ operation independently: calling $backward()$ in PyTorch not only computes the gradient (which is added to any existing gradient), but also frees the computational graph (hence freeing memory). The $step()$ (and $zero\_grad()$) method is then called only once (after six $backward()$ calls. This simple technique reduces the VRAM requirement of training MegaLoc from (roughly) 300GB to 60GB.
@@ 3.2.2 Visual Localization
Image retrieval is a core tool to solve 3D vision tasks, in pipelines like visual localization (e.g. Hierarchical Localization [28] and InLoc [35]) and 3D reconstructions (e.g. COLMAP [31, 30] and GLOMAP [22]). To understand if our method can help this use case, we compute results on the three datasets of Lamar [29], which comprise various challenges, including plenty of visual aliasing from both indoor and outdoor imagery. To do this, we relied on the official LaMAR codebase [^1] https://github.com/microsoft/lamar-benchmark by simply replacing the retrieval method. Results are reported in Tab. 2.

### v3

@@ Abstract
Retrieving images from the same location as a given query is an important component of multiple computer vision tasks, like Visual Place Recognition, Landmark Retrieval, Visual Localization, 3D reconstruction, and SLAM. However, existing solutions are built to specifically work for one of these tasks, and are known to fail when the requirements slightly change or when they meet out-of-distribution data. In this paper we combine a variety of existing methods, training techniques, and datasets to train a retrieval model, called MegaLoc, that is performant on multiple tasks. We find that MegaLoc (1) achieves state of the art on a large number of Visual Place Recognition datasets, (2) impressive results on common Landmark Retrieval datasets, and (3) sets a new state of the art for Visual Localization on the LaMAR datasets, where we only changed the retrieval method to LaMAR’s official localization pipeline. The code for MegaLoc is available at https://github.com/gmberton/MegaLoc
@@ 1 Introduction
Figure 1: Performance comparison of various Visual Place Recognition models. The y axis shows the results averaged across all evaluation datasets, the x axis shows the inference time, and the size of the circle represents the number of parameters per model.
@@
Why does it matter?”, you may ask. Imagine you are doing 3D reconstruction, where image retrieval is a fundamental component, on a collection of diverse scenes (e.g. to create datasets like MegaDepth [18], MegaScenes [37], or for the evergreen Image Matching Challenge [6]). In some cases there would be small scenes (e.g. reconstruction of a fountain), requiring a retrieval model that is able to retrieve nearby images (few meters away), which is something VPR models excel at, but LR models underperform (see [8] Tab. 14). In other cases however, the scene might be large (e.g. a big landmark like a church), with images hundreds of meters away: while LR models are designed for this, VPR models achieve poor results in this situations (see Sec. 3.2.3). Given these considerations, we note how neither VPR nor LR provide models for the diverse cases of 3D reconstructions, creating a gap in literature that is filled by MegaLoc. As another example where a model like MegaLoc is necessary, one can think of Visual Place Recognition (which is also the first step for Visual Localization), where models are evaluated by using a 25 meters threshold (and queries in popular datasets always have at least one positive within 25 meters). However, in the real world the nearest image to a given query might be 100 meters away, and while ideally we would still want to retrieve it, a VPR model is unlikely to work in such case, as it has been trained to ignore anything further away from the camera.
In this paper we demonstrate that, by leveraging a diverse set of data sources and best practices from LR, VPR and VL, we obtain a single image retrieval model that works well across all these tasks.
@@ 2 Method
The core idea of this paper is to fuse data from multiple datasets, and train a single model. We use five datasets containing both outdoor and indoor images (thorough description below) and catering to different image localization tasks: GSV-Cities [1], Mapillary Street-Level Sequences (MSLS) [39], MegaScenes [37], ScanNet [13] and San Francisco eXtra Large (SF-XL) [7]. At each training iteration, we extract six sub-batches of data, one for each dataset (except SF-XL, from which two sub-batches are sampled) and use a multi-similarity loss [38] computed over each sub-batch. Each sub-batch is made of 128 images, containing 4 images (called quadruplets) from 32 different places/classes. Given that these datasets have diverse format, they require different sampling techniques. In the following paragraphs we explain how data is sampled from each dataset.
@@ 3.1 Implementation details
consists of a DINO-v2-base backbone [21] followed by a SALAD [16] aggregation layer, which has shown state-of-the-art performances over multiple VPR datasets [16, 33]. The SALAD layer is computed with 64 clusters, 256 channels per cluster, a global token of 256 and an MLP dimension of 512. The SALAD layer is followed by a linear projection (from a dimension of 16640 to 8448) and an L2 normalization. During training, most of the model is frozen, except for the 4 last transformer layers, the SALAD layer and the final linear layer.
@@
is achieved using PyTorch [23], where instead of summing all the losses and computing a single backward pass, we compute backward passes independently on each dataset: the backward pass computes and accumulates the gradient while freeing memory (see Algorithm 1). This simple technique reduces the VRAM for training MegaLoc from 360GB to 60GB (a 6x reduction because there are 6 losses) with no increase in latency, unlike in standard gradient accumulation where a batch is split into micro batches: this allows MegaLoc to be trained on a single A100 GPU.
Algorithm 1 Memory-Efficient GPU Training
1:Model $Model$, Optim. $Opt$, Datasets $\{D_{i}\}_{i=1}^{N}$
2:Initialize $Model$, $Opt$
3:for each training iteration do
4:     for each dataset $D_{i}$ in $\{D_{1},...,D_{N}\}$ do
5:         $B_{i}\leftarrow\text{LoadBatch}(D_{i})$
6:         $Y_{i}\leftarrow Model(B_{i})$
7:         $L_{i}\leftarrow\text{Loss}(Y_{i},\text{target}_{i})$
8:         $L_{i}.backward()$
9:     end for
10:     $Opt.step()$
11:     $Opt.zero\_grad()$
12:end for
@@ 3.2.2 Visual Localization
Image retrieval is a core tool to solve 3D vision tasks, in pipelines like visual localization (e.g. Hierarchical Localization [28] and InLoc [35]) and 3D reconstructions (e.g. COLMAP [31, 30] and GLOMAP [22]). To understand if our method can help this use case, we compute results on both query sets (phone and HoloLens) of the three datasets of LaMAR [29], which comprise various challenges, including plenty of visual aliasing from both indoor and outdoor imagery. To do this, we relied on the official LaMAR codebase [^1] https://github.com/microsoft/lamar-benchmark by simply replacing the retrieval method. We computed these experiments on the validation set, as the labels for the test set are not publicly available. Note that we only use LaMAR (or the landmark retrieval datasets) as a test set, and using it as validation would take unfeasibly long (multiple days). Results are reported in Tab. 2, and show that while some other models achieve good results on some datasets, MegaLoc is the only one that consistently achieves high results.
@@ 4 Conclusion and limitations
Acknowledgements
Carlo Masone was supported by FAIR - Future Artificial Intelligence Research which received funding from the European Union Next-GenerationEU (Piano nazionale di ripresa e resilienza (PNRR) – missione 4 componente 2, investimento 1.3 – D.D. 1555 11/10/2022, PE00000013). This manuscript reflects only the authors’ views and opinions, neither the European Union nor the European Commission can be considered responsible for them.
We acknowledge the CINECA award under the ISCRA initiative, for the availability of high performance computing resources.